OpenStudy (anonymous):

Which of the following has a solution of 5 3 times the recipricol of 3/5 6÷ 5/6 7/8 of 4 3/8 1/4•1/20

OpenStudy (dangerousjesse):

What's the recipricol of \(\large \frac{3}{5}\)?

OpenStudy (anonymous):

5?

OpenStudy (dangerousjesse):

Are you sure? Here's a hint: \(\huge \frac{1}{\frac{3}{5}}\)

OpenStudy (anonymous):

Question #2 Solve t/2 = 3 2/5

OpenStudy (anonymous):

I don't get it

OpenStudy (dangerousjesse):

Solve \(\large 1\div \frac{3}{5}\) :)

OpenStudy (mathmath333):